OpenStudy (anonymous):

Integrate 1/ cos^2 (x)

OpenStudy (shubhamsrg):

1/cos = sec

OpenStudy (hba):

\[\int\limits_{}^{}1/\cos^2(x) dx =\int\limits_{}^{}\sec^2(x) dx\] Now just integrate

OpenStudy (hba):

@jychay2 Got it ?

OpenStudy (anonymous):

ya,I got it.Thx^^

OpenStudy (anonymous):

This is just a matter of memorizing the derivatives and definitions of certain trigonometric functions.

OpenStudy (hba):

@jychay2 So what should be the answer ?

OpenStudy (anonymous):

tany. Am i right?

OpenStudy (anonymous):

oh,srry.Its tanx

OpenStudy (hba):

No you are wrong the answer would be \[\tan(x)+c\] c is very imp,if you miss it you get no marks+your answer goes wrong.

OpenStudy (hba):

@jychay2 Got it ?

OpenStudy (anonymous):

Oh,ya! Forgot the c,constant. Thx for the correction

OpenStudy (hba):

Your'e welcome :)

OpenStudy (anonymous):

Yeah, saying the answer is \(\tan(x)\) is basically saying that the answer is \(\tan(x) + 0\), which is also like saying the answer is \(\tan(x)+103\). It's like saying the population is \(100,021\) just as \(4\) people enter the city. That's why they are so particular about it. But it is kinda funny too.

OpenStudy (kainui):

Consider that you can derive the result sec^2(x) as the derivative of tanx simply because you know the derivatives of sine and cosine, which is essentially what tangent is. I suggest everyone does this at least once in their life.

OpenStudy (anonymous):

Lol, boring exercise no? \[\frac{d}{dx} \frac{\sin(x)}{\cos(x)}=\frac{\cos^2(x)+\sin^2(x)}{\cos^2(x)}=\sec^2(x)\]